How many grams ofNH

3

N

H

3

 can be produced when 3.60 L

L

 of NO

2

N

O

2

 reacts at 415 

C

C

 and 695 mmHg

m

m

H

g

 ?


1
Expert's answer
2021-10-28T02:59:44-0400

2NO2 + 7H2 → 2NH3 + 4H2O

"V= 3.60 \;L \\

T=415 + 273 = 688 \;K \\

p=695 \;mmHg = 0.914 \;atm"

Ideal gas law

pV =nRT

R = 0.08206 L×atm/mol×K

"n = \frac{pV}{RT} \\

n(NO_2) = \frac{0.914 \times 3.60}{0.08206 \times 688} = 0.0583 \;mol"

According to the reaction:

n(NH3) = n(NO2) = 0.0583 mol

MM(NH3) = 17.0 g/mol

m(NH3) "= 0.0583 \times 17.0 = 0.990 \; g ≈ 1 \;g"

Answer: 1 g